Given the following densities:

Material

Density (kg/m3)
Air 1.28
Ethyl Alcohol 790
Gasoline 680
Water 1000
Glycerin 1260

The bottom of a steel "boat" is a 7 m ✕ 14 m ✕ 2 cm piece of steel (ρsteel = 7900 kg/m3). The sides are made of 1.80 cm thick steel. What minimum height must the sides have for this boat to float in perfectly calm water?

_________ cm
Answer #1

Vs = 7*14*.02 + 2 * .0018 *(14 h + 7 h) = 1.96 + 0.756 h volume of steel
14 * 7 * h = volume of water displaced
14 * 7 * h * 1000 = (1.96 + 0.756 h) *7900 mass of water displaced = mass steel
g cancels out when equating weights
92027.5h = 15484
h = .168 m = 16.8 cm minimum height of sides